# NRR-Phi: Text-to-State Mapping for Ambiguity Preservation in LLM Inference

### v3

@@ Abstract
Implementation: A reference implementation accompanying this work is available at https://github.com/kei-saito-research/nrr-phi.
@@ 5.8 Limitations
Note on LLM role: LLMs serve as convenient sources of human-like interpretation enumeration; the theoretical claims of this paper concern the structure of the resulting states, not the capabilities of the underlying models. The $\phi$ mapping framework is model-agnostic and could be instantiated with any interpretation source.
@@ 8 Conclusion
Experimental evaluation on 68 ambiguous sentences demonstrates that the mapping preserves interpretive multiplicity: mean state entropy $H=1.087$ bits across all categories, compared to $H=0$ for collapse-based baselines.
@@ D.2.1 Core Transition Operators (Empirically Validated)
This operation multiplies all weights by the same factor $\left(1-b/\max_{j}w_{j}\right)$. Therefore, after normalization for entropy evaluation, the induced probability distribution is unchanged and $\Delta H=0$. We retain this as a scale-equivariance calibration operator, not as an entropy-changing update rule. This $\sigma_{\mathrm{state}}$ is separate from item-level strengthen operators used in NRR-IME/Universal.
@@ D.6.3 Key Findings
    Dampening increases entropy: $\delta$ v2 shifts distributions toward uniformity, raising $H$ by an average of 0.03 bits at $\lambda=0.3$. This increase is analytically guaranteed by the strict concavity of Shannon entropy (see Appendix D) and confirmed empirically.

### v4

@@ Abstract
Implementation: A reference implementation accompanying this work is available at https://github.com/kei-saito-research/nrr-phi. Series hub: A cross-paper index for the NRR series is available at https://github.com/kei-saito-research/nrr-series-hub.
@@ 1.4 Paper Organization
1.5 Series Alignment: Notation and Metrics
Notation contract. In this paper, $\sigma_{\mathrm{state}}$ denotes a state-level calibration operator in Appendix D. This is distinct from the item-level strengthen/dampen operators ($\sigma_{\text{item}},\delta_{\text{item}}$) used in NRR-IME/Transfer. The symbol family is consistent; the operation level differs.
Metric contract. Phi primarily evaluates text-to-state validity and non-collapse properties (state entropy and violation rates). Later implementation papers add interface/operational metrics such as extraction reliability and token efficiency. Cross-paper interpretation should treat these as layered metrics across stages, not as competing definitions of success.
Boundary statement. Claims are limited to tested datasets and operator validation settings in this paper. In this paper, validity is condition-indexed rather than unconditional. A mapped interpretation is valid under its associated context tag, not universally valid across contexts.
@@ 5.8 Limitations
Note on LLM role: LLMs serve as convenient sources of human-like interpretation enumeration; the theoretical claims of this paper concern the structure of the resulting states, not the capabilities of the underlying models. The $\phi$ mapping framework is model-agnostic and could be instantiated with any interpretation source. NRR is a derivation-and-evaluation framework, not a closed recipe; the same derivation lens supports alternative mapping/operator instantiations under fixed protocols with condition-bounded reporting.
@@ 7.1 Implications for LLM Architecture
In this framing, collapse and commitment are distinct: collapse is a point event at the output boundary, whereas commit inertia is trajectory-level persistence of prior selections across turns. The $\phi$ mapping and non-destructive projection reduce this inertia by keeping alternatives available after each output step. This frame-first design is intentionally structural: by preserving parallel candidate frames after each output step, systems reduce the need for unilateral early forcing in subsequent turns. Accordingly, structural readings are frame-conditional in operation: the same textual surface can map to different relation-structures under different active frame sets, without logical contradiction. We treat this as a condition-bounded architectural implication that downstream papers should test empirically, rather than as a universal interpersonal claim.
@@ 7.4 Future Work
Reproducibility note.
We release code, fixed settings, and primary logs for the reported mapping and operator-validation runs. The executable commands fix the processing mode, input files, and output targets, with no temperature-based sampling in these offline scripts. Model-related prompt artifacts are published as repository snapshots, and deterministic rule-based outputs are regenerated from the same scripts. The full environment specification and artifact map are provided in nrr-phi/reproducibility.md.
@@ 8 Conclusion
Experimental evaluation on 68 ambiguous sentences demonstrates that the mapping preserves interpretive multiplicity: mean state entropy $H=1.087$ bits across all categories, compared to $H=0$ for collapse-based baselines. This does not reject collapse itself; it targets premature collapse and preserves explicit commitment at the output boundary when conditions support reliable resolution. In this setup, improvements come from reducing unnecessary early commitment and preserving reversible update paths, not from adding architectural complexity. Series Consistency Statement: Unsupported interpretations are never injected at a fixed evidence state. Hypothesis-set expansion is evidence-gated, and unobserved possibilities remain unresolved. Evidence in this paper is restricted to explicit input, provided context, and declared retrieval outputs.
@@ D.2.1 Core Transition Operators (Empirically Validated)
This operation multiplies all weights by the same factor $\left(1-b/\max_{j}w_{j}\right)$. Therefore, after normalization for entropy evaluation, the induced probability distribution is unchanged and $\Delta H=0$. We retain this as a scale-equivariance calibration operator, not as an entropy-changing update rule. This $\sigma_{\mathrm{state}}$ is separate from item-level strengthen operators used in NRR-IME/Transfer.
@@ D.6.3 Key Findings
    Dampening increases entropy: $\delta$ v2 shifts distributions toward uniformity, raising $H$ by an average of 0.03 bits at $\lambda=0.3$. This increase is analytically guaranteed by the strict concavity of Shannon entropy (see Appendix D) and confirmed empirically. Here, entropy increase denotes redistribution across already extracted interpretations, not insertion of unsupported new interpretations.